How are the digits in a number grouped?
In much of the world, in groups of three from the decimal point outwards (in both directions). On the integer side the groups may be separated by commas (or full stops).
In South Asia, though, integers are grouped in three and then in twos - again moving to the left from the decimal point.
What is a grouped frequency distribution used for?
What is a independant and dependant variale?
An independent variable is the factor that is manipulated or changed in an experiment to observe its effects, while the dependent variable is the outcome or response that is measured as a result of the changes made to the independent variable. For example, in a study examining the effect of fertilizer on plant growth, the amount of fertilizer used is the independent variable, and the growth of the plants is the dependent variable. The relationship between the two helps researchers understand cause-and-effect dynamics.
Where can you find reloading data for Remington componets?
Common sources include the loading manuals from Speer, Lyman, Hornaday
Loading manuals for companies like Speer and Hornady are only going to give you load data for THEIR bullets--not Remington bullets. I haven't checked Lyman yet but the load data published by powder manufacturers are typically going to list loads using bullets manufactured by the major bullet companies. I am looking for a source that publishes loads using Remington bullets myself. No, you can't simply substitute a load using a bullet of the same weight as the Remington bullet you want to use. Bullet composition and design impacts CUP's.
1. Better chance of uniform sample.
2. Material for confirmations if needed.
How many diamonds are in a standard deck of cards?
There are 13 diamonds in a standard deck of 52 cards.
The largest city in the world is Tokyo, Japan (population: 32,450,000), followed by Seoul, South Korea (20,550,000) and Mexico City, Mexico (20,450,000).
99
What are the odds of rolling 3-6's or 8's before rolling a 7 in craps?
Approximately 1/1120 * You have a 27.777% chance of rolling either a six or an eight on the first roll. * You have a 13.888% chance of rolling the same thing as the first roll. * You have a 13.888% chance of doing it again. * You have a 16.666% chance of rolling a seven. To find total probability in this situation, you multiply all of your chances together. .27777 x .13888 x .13888 x .16666 = 0.000892886 = .0892886% =1/1119.964 You may just be finding this out now, but casinos have known it for a long time. The casinos simply make their payouts less than the probability and they can't lose. While I recognize that your question deals with the probablity of an event recurring, for the average gambler, there is another way to look at this. Dice have no memory. Therefore, each time you roll them, the probability of rolling a given number remains the same. Each time you roll the dice, the chances of rolling a 6 before a 7 is 6 to 5 against you. The same odds apply to rolling an 8 before a 7.
What does the term Prima Donna mean in Opera?
Literally, it means "First Lady" in Italian.
The term has also come to refer to a conceited, temperamental individual.
What are the multiples of three?
3 6 9 12 15 18 21 24 27 30 33 36 39 42 45 48 51 54 57 60 63 66 69 72
What types od individuals in a population are represented by the two ends of a bell curve?
The individuals with extreme variations of a trait.
The rate at which population size grows or declines depends on the rate of what?
births, death, immigration and emigration
How do you interpret thermatic data?
Bung all the data into Excel and download the Thermatic Data Analysis Package DLL from Microsoft. Worked for me.